I'm having trouble with fractions as exponents. Simplify the expression, if possible.
a) 81^3/4
b) -81 3/4
c) (-81)^3/4
d) 81^-3/4
e) -81^-3/4
f) (-81)^-3/4
Please help!

Answer by checkley71(8403)

A) 81^3/4=THE FOURTH ROOT OF 81^3 OR FOURTH ROOT OF (81*81*81) OR
FOURTH ROOT OF (9*9*9*9*9*9) OR 9FOURTH ROOT OF (9*9)
OR 9FOURTH ROOT(3*3*3*3)
OR 9*3=27 ANSWER.
PROOF
27^4=81^3
531,441=531,441.